Ardin:

Im unsure what it wants in the blanks, i remember sine and cosine now but, does it want the formula plus the formula? I don't know why it would...

Ardin:

1 attachment
Ardin:

Im just very confused since normally i'm supposed to find the numbers but the only number here is the 90 degrees i just, ive never had to put the formula in the blanks that way its weird

snowflake0531:

\[sin(\theta)=\frac{opposite}{hypotenuse}\\cos(\theta)=\frac{adjacent}{hypotenuse}\\tan(\theta)=\frac{opposite}{adjacent}\]

Ardin:

@extrinix wrote:
Ardin, just a question, do you want me to answer the one that you closed earlier?
oh, no your good, i figured it out soon after posting it

Ardin:

im not sure what to put for this tbh

snowflake0531:

Well, first tell me, what's sin B and tan C

Ardin:

it doesn't have a number, its just b/a and c/b

Ardin:

if i did that right

snowflake0531:

Yes so what is \[\frac{b}{a} \times \frac{c}{b}\]

Ardin:

ohh, uh should be bc/ab right?

snowflake0531:

yes but the b's cancel out

Ardin:

so c/a?

snowflake0531:

yeah

snowflake0531:

and what's sinC and tan B?

Ardin:

c/a and b/c should be

snowflake0531:

\[\frac{ c }{ a } \times \frac{ b }{ c }\] the c's can cancel out

Ardin:

b/a?

snowflake0531:

yes

Ardin:

ok, thank you!

snowflake0531:

yw~

Extrinix:

 Just a guide if you ever need it.

Ardin:

@extrinix wrote:
Created with RaphaƫlabcBACadjacent of Aopposite of Bhypotenuse of A B Copposite of C, adjacent of BReply Using Drawing Just a guide if you ever need it.
Thank you!